find an equation of the circle centered at (6,-5) and passing through (1,7)
thank you

Answer by stanbon(48558)

find an equation of the circle centered at (6,-5) and passing through (1,7)
thank you
------------
If it passes thru (1,7) the length of the radius is:
r=sqrt[(7--5)^2+(1-6)^2]=sqrt[144+25]=sqrt169=13
----------
With the center at (6.-5) and radius = 13 the equation is:
(x-6)^2 + (y+5)^2=13^2
---------
Cheers,
Stan H.